OpenStudy (anonymous):

Help with algebra for Medal and Fan! Select the proper equation in point-slope form for the given point and slope. Point: (–2, 4); Slope: 5 (Points : 1) A. y + 4 = 5x – 10 B. y – 4 = 5(x – 2) C. y + 4 = 5(x + 2) D. y – 4 = 5(x + 2)

OpenStudy (anonymous):

ok exactly the same as the last one precisely

Nnesha (nnesha):

point slope intercept form is \[y - y _{1} = m ( x - x _{1})\] slope is 5 so just replace that by m

OpenStudy (anonymous):

only the numbers are different

Nnesha (nnesha):

use this point ( -2 , 4) for \[(x _{1} , y _{1})\]

OpenStudy (anonymous):

Point: (-2, 4); Slope: 5 ............x1,y1........... m C.y - y1 = m(x - x1) y - 4 = 5(x - (-2)) y - 4 = 5(x + 2) Point: (1, -7); Slope: -2/3 A. y + 7 = -2/3 (x - 1) Point: (5, -9); Slope: 2 C. y + 9 = 2(x - 5)

OpenStudy (anonymous):

\[\huge y-\color{red}{y_1}=\color{green}m(x-\color{blue}{x_1})\]

OpenStudy (anonymous):

y-4=5(x+2) is that right @satellite73?

Nnesha (nnesha):

yesh

OpenStudy (anonymous):

so who do i give the medal to and who do i fan?